I need to buy tires. Standard set $600 last 3 years or $800 last 4 years. Which do you buy? With my calculations I believe it is the $800 tires with only a difference of $96.
1
Expert's answer
2020-04-29T09:22:45-0400

If $600 set lasts 3 years and $800 set lasts 4 years, then the average cost per year of usage is 600/3 = 800/4 = 200 in both cases. So, no matter what tires will you buy.